OpenStudy (anonymous):

The area of a triangle is 24 square inches. What is the height of the triangle if the base length is 6 inches?

OpenStudy (anonymous):

8in?

OpenStudy (anonymous):

@mathmath333 can you help

OpenStudy (mathmath333):

use the formula area of triangle=\(\dfrac{1}{2}\times height \times base\)

OpenStudy (mathmath333):

8 in is correct

OpenStudy (anonymous):

@mathmath333 thanks